A 13.97 g sample of ice is heated from -10 oC to its melting point of 0 oC and then melted completely. Next it is raised to its boiling point of 100 oC and boiled completely. Finally the sample is heated to 111 oC. How much energy was absorbed by the sample? (Answer in kJ pay close attention to units!)

Sice=2.1 J/goC Delta Hf = 6.02 kJ/mol

Ssteam=1.7 J/goC Delta Hv = 2.26 kJ/g


1
Expert's answer
2021-03-23T04:22:00-0400

Solution:

Moles of H2O = Mass of H2O / Molar mass of H2O

The molar mass of H2O is 18.0153 g mol-1.

Hence,

Moles of H2O = 13.97 g / 18.0153 g mol-1 = 0.77545 mol

Moles of H2O = 0.77545 mol


This problem can be summarized thusly:

1) Ice is heated from -10°C to 0°C.

The heat absorbed is calculated by using the specific heat of ice (Sice = 2.1 J/g°C) and the equation:

q = m × S × ΔT

q1 = m × Sice × (Tf - Ti)

q1 = 13.97 g × 2.1 J/g°C × (0 - (-10))°C = 293.37 J = 0.29337 kJ

q1 = 0.29337 kJ


2) Ice is melted at 0°C. The heat absorbed is calculated by multiplying the moles of ice by the molar heat of fusion (ΔHf).

q2 = ΔHf × n = 6.02 kJ/mol × 0.77545 mol = 4.66821 kJ

q2 = 4.66821 kJ


3) Water at 0°C is heated to 100°C. The heat absorbed is calculated by using the specific heat of water (Swater = 4.2 J/g°C) and the equation: q = m × S × ΔT

q3 = m × Swater × (Tf - Ti)

q3 = 13.97 g × 4.2 J/g°C × (100 - 0)°C = 5867.4 J = 5.8674 kJ

q3 = 5.8674 kJ


4) Water is vaporized to steam at 100°C. The heat absorbed is calculated by multiplying the grams of water by the heat of vaporization (ΔHv) in kJ/g.

q4 = ΔHv × m = 2.26 kJ/g × 13.97 g = 31.5722 kJ

q4 = 31.5722 kJ


5) Steam is heated from 100°C to 111°C. The heat absorbed is calculated by using the specific heat of steam (Ssteam = 1.7 J/g°C) and the equation: q = m × S × ΔT

q5 = m × Ssteam × (Tf - Ti)

q5 = 13.97 g × 1.7 J/g°C × (111 - 100)°C = 261.239 J = 0.26124 kJ

q5 = 0.26124 kJ


q = q1 + q2 + q3 + q4 + q5

q = 0.29337 kJ + 4.66821 kJ + 5.8674 kJ + 31.5722 kJ + 0.26124 kJ = 42.66242 kJ = 42.66 kJ

q = 42.66 kJ


Answer: 42.66 kJ of energy was absorbed by the sample.